# Effect of the scalar unparticle and polarization at muon colliders through the exclusive W boson hadronic decays in the final state in the Randall-Sundrum model

### v3

@@ Abstract
In this paper, by using Feynman diagram techniques we have evaluated the effect of the scalar unparticle and polarization at the muon colliders through the exclusive W boson hadronic decays $W^{\pm}\rightarrow\pi^{\pm}\gamma$, $W^{\pm}\rightarrow K^{\pm}\gamma$, $W^{\pm}\rightarrow\rho^{\pm}\gamma$ in the Randall-Sundrum (RS) model. The results show that with fixed collision energies, the total cross-sections for hadronic productions in final states depend strongly on the parameters of the unparticle physics and muon beam polarization. The total cross-sections achieve the maximum value when both of muon beams polarize left or right and the minimum value when the $\mu^{-}$ beam polarizes left, the $\mu^{+}$ beam polarizes right and vice versa. In case of the different polarization, the cross-sections increase as the collision energy increases and they change insignificantly when the $\mu^{-}$ beam polarizes left, the $\mu^{+}$ beam polarizes right. With the benchmark background $(\Lambda_{U},d_{U})$ =(1=(1= ( 1TeV,1.9),1.9), 1.9 ), the cross-sections reach the maximum value. The integrated luminosity value is shown to correspond to a significance $1\sigma,3\sigma$.
@@ I Introduction
After the discovery of the Higgs particle at the Large Hadron Collider (LHC) at CERN [1, 2], the particle content of the Standard Model (SM) has finally been completed. However, there are many open questions in physics that SM has yet to be address. One of the most attractive extended models beyond the SM is the Randall-Sundrum (RS) model [3]. The existence of an additional scalar called the radion ($\phi$) is corresponded to the quantum fluctuations of the distance between the two three-branes: UV brane and IR brane. Radion and Higgs boson have the same quantum numbers. General covariance allows a possibility of mixing between the radion and the Higgs boson [4, 5]. In the Lagrangian of SM, the scale invariance is broken at or above the electroweak scale [6, 7]. At TeV scale, the scale invariant sector has been considered as an effective theory and that if it exists, it is made of unparticle suggested by Geogri [8, 9] and may become part of reality. Based on the Banks-Zaks theory [10, 11], unparticle stuff with nontrivial scaling dimension is considered to exist in our world and this opens a window to test the effects of the possible scalar invariant sector, experimentally. The effects of unparticle on properties of high energy colliders have been intensively studied in Refs.[12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23]. Dark matter and unparticles production in association with a Z boson in pp collisions at the center-of-mass energy 8 TeV has been considered in Ref.[24]. Recently, the scalar unparticle signals at LHC are studied in detail in Ref.[25]. The anomalous couplings at LHeC is researched in Ref.[26]. The muon colliders, which can reach center-of-mass energy up to tens of TeV, provide an unprecedented potential in probing new physics beyond the Standard Model[27]. The advantage of initial muon beam polarization is that it is effective for the indirect search [28, 29]. It is versatile in the fundamental particle physics and nuclear physics [30, 31]. The integrated luminosity scaling of a high energy muon collider (assuming a 5 year run) reachs 1 $ab^{-1}$ (3 TeV), 10 $ab^{-1}$ (10 TeV) and 20 $ab^{-1}$ (14 TeV) [32, 33, 34, 35]. In our previous work, investigation of the scalar unparticle and anomalous couplings at muon colliders in final states with multiple photons in the RS model is considered in detail in Ref.[36]. The contribution of the scalar unparticle on the WW production at ILC is studied in Ref.[37] in which we only considered in s-channel without considering the influence of $\gamma WW,ZWW$ anomalous couplings and neutrino propagator. Moreover, rare hadronic decays of W bosons in the final states would provide an accurate measurement of the W boson mass that is based solely on visible decay products at future colliders [38, 39, 40]. These rare processes, if observed, would validate the quantum chromodynamics (QCD) factorization formalism used to calculate cross-sections at colliders [40].
@@
          In this work, by using the rare decay channels of W boson $W^{\pm}\rightarrow\pi^{\pm}\gamma$, $W^{\pm}\rightarrow K^{\pm}\gamma$, $W^{\pm}\rightarrow\rho^{\pm}\gamma$ we calculate in detail the cross-sections for hardronic production in the final states at muon colliders in the RS model. This paper is arranged as follows. The effect of the scalar unparticle and polarization at the muon colliders through the exclusive W boson hadronic decays are calculated in detail in Section II. Finally, we summarize our results and make conclusions in Section III.
@@
II Effect of the scalar unparticle and polarization at the muon colliders through the exclusive W boson hadronic decays
@@
Influence of unparticle and polarization on properties of high energy colliders have been intensively studied in Refs.[12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23]. The production of dark matter and unparticles associated with Z boson at 8 TeV via pp collisions has been analyzed in Ref.[24]. Recently, the scalar unparticle signals at LHC 14 TeV are investigated in Ref.[25]. Phenomenology of heavy neutral gauge boson at muon colliders are peresented in detail in [41], which indicate that it can provide an unprecedented potential in probing new physics beyond the Standard Model.
@@
          Study for the exclusive boson hadronic decay with the ATLAS detector at 13 TeV is considered in detail in Ref.[40]. Recently, mass biases in reconstructing exclusive radiative hadronic decays of W boson at the LHC are given in Ref.[42]. Rare and exclusive few-body decays of the Higgs, Z and W bosons and the top quark are given in Ref.[43]. However, the influence of scalar unparticle and polarization at muon colliders through hadronic decays of W boson has not yet been invested. In this work, by using the rare decay channels of W boson $W^{\pm}\rightarrow\pi^{\pm}\gamma$, $W^{\pm}\rightarrow K^{\pm}\gamma$, $W^{\pm}\rightarrow\rho^{\pm}\gamma$ we calculate in detail the cross-sections for hardronic production at muon colliders.
@@
$M_{Z}=-i\dfrac{g}{2cos\theta_{W}(q^{2}_{s}-m^{2}_{Z})}\overline{v}(p_{2})\gamma^{\sigma}\left(-\frac{1}{2}+2sin^{2}\theta_{W}-\frac{1}{2}\gamma_{5}\right)u(p_{1})\left(\eta_{\sigma\beta}-\dfrac{q_{s\sigma}q_{s\beta}}{m^{2}_{Z}}\right)\varepsilon^{*}_{\mu}(k_{1})\Gamma_{WWZ}^{\beta\mu\nu}\varepsilon^{*}_{\nu}(k_{2})$ (4)
@@
$M_{U}=i\overline{g}_{\mu\mu U}\overline{g}_{WWU}\dfrac{A_{d_{U}}}{2sin(d_{U}\pi)}(-q^{2}_{s})^{d_{U}-2}\overline{v}(p_{2})u(p_{1})\varepsilon^{*}_{\mu}(k_{1})\left[\left(k_{1}k_{2}\right)\eta^{\mu\nu}-k_{1}^{\nu}k_{2}^{\mu}\right]\varepsilon^{*}_{\nu}(k_{2}),$ (7)
Here, $q_{s}=p_{1}+p_{2}=k_{1}+k_{2}$, $s=(p_{1}+p_{2})^{2}$ is the square of the collision energy, $M_{U}$ is the contribution by the scalar unparticle, which is important for the described process. $\overline{g}_{\mu\mu\phi},\overline{g}_{\mu\mu h},\overline{g}_{W\phi},\overline{g}_{Wh}$ are the couplings of Higgs/radion shown in Ref.[4]. $\overline{g}_{\mu\mu U},\overline{g}_{WWU},A_{d_{U}}$ are given by Ref.[7, 9]. Anomalous couplings are shown in Ref.[44] as follow
@@
$M_{t}=-\frac{g^{2}}{2}\varepsilon^{*}_{\nu}(k_{2})\overline{v}(p_{2})\gamma^{\mu}\frac{1-\gamma^{5}}{2}\widehat{q}_{t}\varepsilon^{*}_{\mu}(k_{1})\gamma^{\nu}\frac{1-\gamma^{5}}{2}u(p_{1}).$ (11)
The total cross-section for the whole process can be calculated as follow
@@
where $\psi=(\overrightarrow{p}_{1},\overrightarrow{k}_{1})$ is the scattering angle. The model parameters are chosen as $\lambda_{\mu\mu}=\lambda_{WW}=\lambda_{0}=1$, $m_{h}$ = 125 GeV, $m_{\phi}$ = 125 GeV, $\Lambda_{\phi}=5$ TeV [26]. The integrated luminosity scaling of a high energy muon collider reachs 140 $fb^{-1}$ at a center-of-mass energy of $\sqrt{s}=13$ TeV [40]. The bounds on the anomalous $W^{-}W^{+}\gamma$ and $W^{-}W^{+}Z$ couplings are provided by the LEP, Tevatron and LHC experiments. The ALTAS collaboration has updated the best available constraints on anomalous couplings $\Delta k_{\gamma}$, $\lambda_{\gamma}$, $\Delta k_{Z}$, $\lambda_{Z}$ obtained as follows: $\Delta k_{\gamma}\in[-0.135,0.190]$, $\lambda_{\gamma}\in[-0.065,0.061]$, $\Delta k_{Z}\in[-0.061,0.093]$, $\lambda_{Z}\in[-0.062,0.065]$ [46]. From the formulas (28 - 30) above and branching ratio of W boson in Ref.[40], we give estimates in detail for the cross-sections as follows:
      i) In Fig.1, the total cross-sections are plotted as the function of $P_{\mu^{-}},P_{\mu^{+}}$, which are the polarization coefficients of $\mu^{-},\mu^{+}$ beams, respectively. The parameters are chosen as $d_{U}=1.1$, $\Lambda_{U}=1$ TeV [36], $\sqrt{s}=13$ TeV. The figures indicate that the total cross-sections achieve the maximum values when $P_{\mu^{-}}=P_{\mu^{+}}=\pm 1$ and the minimum values when $P_{\mu^{-}}=1,P_{\mu^{+}}=-1$ or $P_{\mu^{-}}=-1,P_{\mu^{+}}=1$.
      ii) The model parameters are chosen as in Fig.1. The polarization coefficients $(P_{\mu^{-}},P_{\mu^{+}})$ are chosen as in Ref. [41], $(P_{\mu^{-}},P_{\mu^{+}})$ $=(1,-1),(0.8,-0.8),(0.6,-0.6),(1,1)$, respectively. The total cross-sections in $\mu^{+}\mu^{-}\rightarrow W^{+}W^{-}\rightarrow\pi^{-}\pi^{+}\gamma\gamma/K^{-}K^{+}\gamma\gamma/\rho^{-}\rho^{+}\gamma\gamma$ collisions which depend on the collision energy are shown in Fig.2. From the figure we can see that the cross-sections increase when the collision energy increases. In case of $\mu^{-}$ beam is left polarized, $\mu^{+}$ beam is right polarized and vice versa, the cross-sections change insignificantly.
      iii) The total cross - section depends on ($\Delta k_{\gamma}$, $\lambda_{\gamma}$) shown in the Fig.3. The parameters are chosen as $P_{\mu^{-}}=0.8,P_{\mu^{+}}=-0.8$, $\sqrt{s}=13$ TeV, $\Lambda_{U}=1$ TeV, $d_{U}=1.1$, $\Delta k_{Z}=0.093$, $\lambda_{Z}=0.065$. With the fixed value of $\Delta k_{\gamma}$, the cross-sections in $\mu^{+}\mu^{-}\rightarrow W^{+}W^{-}\rightarrow\pi^{-}\pi^{+}\gamma\gamma/K^{-}K^{+}\gamma\gamma/\rho^{-}\rho^{+}\gamma\gamma$ collisions are independent on $\lambda_{\gamma}$ values. The result shows that cross-sections are the largest in the yellow region of the figures in which numerical values are given by (a) $2.9\times 10^{-4}$ fb, (b) $2.35\times 10^{-4}$ fb and (c) $2.2\times 10^{-3}$ fb in final states, respectively.
      iv) The total cross - section depends on ($\Delta k_{Z}$, $\lambda_{Z}$) shown in the Fig.4. The model parameters are chosen as in Fig.3, $P_{\mu^{-}}=0.8,P_{\mu^{+}}=-0.8$, $\sqrt{s}=13$ TeV, $\Lambda_{U}=1$ TeV, $d_{U}=1.1$, $\Delta k_{\gamma}=0.190$, $\lambda_{\gamma}=0.061$. The cross-sections in the yellow region of the figures are the largest with the values (a) $2.995\times 10^{-4}$ fb, (b) $2.395\times 10^{-4}$ fb and (c) $2.245\times 10^{-3}$ fb in final states, respectively.

### v5

# The influence of the scalar unparticle and polarization on the exclusive W boson hadronic decays in the final state at muon colliders in the Randall-Sundrum model

@@ Abstract
An attempt is made to present the effect of unparticle physics and polarization on the exclusive decays of W boson at high energy colliders in the Randall-Sundrum (RS) model. By using Feynman diagram techniques we have evaluated the influence of the scalar unparticle and polarization on the exclusive W boson hadronic decays of $W^{\pm}\rightarrow\pi^{\pm}\gamma$, $W^{\pm}\rightarrow K^{\pm}\gamma$ and $W^{\pm}\rightarrow\rho^{\pm}\gamma$ at the high energy muon colliders in the RS model. The result shows that with fixed collision energies, the total cross-section for hadronic productions in the final state depends strongly on the parameters of the unparticle physics and muon beam polarizes. With a center-of-mass energy of 10 TeV, the total cross-sections achieve the maximum value when the benchmark signal point as $(\Lambda_{U},d_{U})$ $(1\text{TeV},1.9)$ and the polarization coefficient as $(P_{\mu^{-}},P_{\mu^{+}})=(1,1)$. The numerical evaluation for the statistical significance is given in detail which indicates that the effect is greatly enhanced in the future experiments with moderately high energy colliders.
@@ I Introduction
One of the most attractive extended models beyond the Standard model (SM) is the Randall-Sundrum (RS) model to solve the hierarchy problem naturally [1]. The RS model setup involves two 3-branes bounding a slice of 5D compact anti-de Sitter space. Gravity is localized UV brane, while the SM fields are supposed to be localized IR brane. The separation between the two 3-branes leads directly to the existence of an additional scalar called the radion ($\phi$), corresponding to the quantum fluctuations of the distance between the two 3-branes [2, 3, 4]. Radion and Higgs boson have the same quantum numbers, the general covariance allows a possibility of mixing between the radion and the Higgs boson [4, 5]. Phenomenology of scalar particles (Higgs/radion) at the high energy colliders in the RS model has been intensively studied in Refs.[6, 7, 8, 9, 10, 11, 12].
     At TeV scale, the scale invariant sector has been considered as an effective theory and that if it exists, it is made of unparticle suggested by Geogri [13, 14] and may become part of reality. The very high energy theory contains the fields of the SM and the fields of a theory with a nontrivial IR fixed point which is called Banks-Zaks (BZ) fields [15]. These have the generic form between the SM and BZ fields as follows [16, 17].
$\dfrac{1}{M_{U}^{k}}\mathcal{O}_{SM}\mathcal{O}_{BZ},$ (1)
where $\mathcal{O}_{SM}$ is an operator with mass dimension $d_{SM}$ built out of SM fields and $\mathcal{O}_{BZ}$ is an operator with mass dimension $d_{BZ}$ built out of BZ fields. In the effective theory below the scale $\Lambda_{U}$, the BZ operators match onto unparticle operators and the interactions of (1) match onto interactions of the form
$\dfrac{C_{U}\Lambda_{U}^{d_{BZ}-d_{U}}}{M_{U}^{k}}\mathcal{O}_{SM}\mathcal{O}_{U},$ (2)
here $d_{U}$ is the scaling dimension of the unparticle operator $\mathcal{O}_{U}$. Unitarity constraints derived from Conformal Field Theory (CFT) set the fundamental lower bound on the dimension at $d_{U}\geq 1$. However, operators with dimensions $d_{U}\leq 2$ are preferentially considered, exceeding this threshold ($d_{U}\geq 2$) makes the calculations unpredictable due to an increased sensitivity to the ultraviolet sector. Therefore, the range $1<d_{U}<2$ is considered [18]. Based on the BZ theory [15, 17], unparticle stuff with nontrivial scaling dimension is considered to exist in our world and this opens a window to test the effects of the possible scalar invariant sector, experimentally. The effects of unparticle on properties of high energy colliders have been intensively studied in Refs.[19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30]. Dark matter and unparticles production in association with a Z boson in pp collisions at the center-of-mass energy 8 TeV has been considered in Ref.[31]. Recently, the scalar unparticle signals at LHC are studied in detail in Ref.[32]. The anomalous couplings at LHeC is researched in Ref.[33].
@@
     The muon colliders, which can reach center-of-mass energy up to tens of TeV, provide an unprecedented potential in probing new physics beyond the SM[34]. The advantage of initial muon beam polarization is that it is effective for the indirect search [35, 36], which is versatile in the fundamental particle physics and nuclear physics [37]. The integrated luminosity scaling of a high energy muon collider (assuming a 5-year run) reaches 1 $\text{ab}^{-1}$ (3 TeV), 10 $\text{ab}^{-1}$ (10 TeV) and 20 $\text{ab}^{-1}$ (14 TeV)[38, 39, 40, 41]. Given that the radion mixes with the Higgs boson, precision measurements of Higgs properties at muon colliders will provide crucial information [42, 43]. In our previous work, investigation of the scalar unparticle and anomalous couplings at muon colliders in final states with multiple photons in the RS model is considered in detail in Ref.[44]. The contribution of the scalar unparticle on the WW production at ILC is studied in Ref.[45] in which we only considered in s-channel without considering the influence of $\gamma WW,ZWW$ anomalous couplings and neutrino propagator. Moreover, the rare hadronic decays of W bosons in the final states would provide an accurate measurement of the W boson mass that is based solely on visible decay products at future colliders [46, 47, 48]. These rare processes, if observed, would validate the quantum chromodynamics (QCD) factorization formalism used to calculate cross-sections at colliders [48].
@@
     In this work, by using the rare decay channels of $W^{\pm}\rightarrow\pi^{\pm}\gamma$, $W^{\pm}\rightarrow K^{\pm}\gamma$, $W^{\pm}\rightarrow\rho^{\pm}\gamma$, we evaluate the influence of the scalar unparticle and polarization on the hadronic production at muon colliders in the RS model. This paper is arranged as follows. The theoretical framework is introduced in Section II. The influence of the scalar unparticle and polarization on the exclusive W boson hadronic decays in the final state at muon colliders are calculated in detail in Section III. Finally, we summarize our results and make conclusions in Section IV.
II Theoretical framework of the Higgs - radion mixing in Randall-Sundrum model
The RS model is based on a 5D spacetime with two three-branes: the UV brane and the IR brane. All the SM and dark matter (DM) fields, excepted for gravity, are assumed at IR brane. Gravity lives on the second three-branes. The RS model consider a non-factorizable 5-dimensional metric in the form [6]
$d_{s}^{2}=e^{-2\sigma}\eta_{\mu\nu}dx^{\mu}dx^{\nu}-r_{c}^{2}dy^{2},$ (3)
where $\sigma=kr_{c}|y|$, $k$ is the curvature along the 5th-dimension, $r_{c}$ is the length-scale which is related to the size of the extra-dimension. For $\sigma=kr_{c}\sim 10$ the RS scenario can address the hierarchy problem. The action in $5\mathrm{D}$ is
$S=S_{\text{gravity }}+S_{\mathrm{IR}}+S_{\mathrm{UV}}.$ (4)
where
$S_{gravity}=\dfrac{16\pi}{M_{5}^{3}}\int d^{4}x\int^{\pi}_{0}r_{c}\sqrt{G^{(5)}}[R^{(5)}-2\Lambda_{5}]dy,$ (5)
$S_{\mathrm{IR}},S_{\mathrm{UV}}$ are the brane actions for the two 3-branes as follows
$S_{\mathrm{IR}}=\int d^{4}x\sqrt{-g}\left(-f^{4}_{IR}+\mathcal{L}_{SM}+\mathcal{L}_{DM}\right),$ (6)
[…]
$M_{5}$ is the fundamental gravitational scale, $G^{(5)}_{MN}$ is the 5-dimensional metric, $R^{(5)}$ is the Ricci scalar, $\Lambda_{5}$ is the 5-dimensional cosmological constant, $f_{IR},f_{UV}$ are the brane tensions for the two 3-branes and $f^{4}_{IR}=-f^{4}_{UV}=\sqrt{-24M^{3}_{5}\Lambda_{5}}$. On the IR brane, the effective gravitational coupling is $\Lambda=\overline{M}_{P}e^{-k\pi r_{c}}$. The 4-dimensional component of the metric at first order is expanded
$G^{(5)}_{MN}=e^{-2\sigma}(\eta_{\mu\nu}+\kappa_{5}h_{\mu\nu}),$ (8)
where $\kappa_{5}=2M_{5}^{-2/3}$, the 5-dimensional field $h_{\mu\nu}$ can be written $h_{\mu\nu}(x,y)=\sum h_{\mu\nu}^{n}(x)\dfrac{\chi^{n}(y)}{\sqrt{r_{c}}}$.
     The energy-momentum tensor is given by
$T_{\mu\nu}=T_{\mu\nu}^{SM}+T_{\mu\nu}^{DM},$ (9)
where
$T_{\mu\nu}^{SM}=$ ${\left[\frac{i}{4}\bar{\psi}\left(\gamma_{\mu}D_{\nu}+\gamma_{\nu}D_{\mu}\right)\psi-\frac{i}{4}\left(\gamma_{\mu}D_{\nu}\bar{\psi}\gamma_{\mu}+D_{\mu}\bar{\psi}\gamma_{\nu}\right)\psi-\eta_{\mu\nu}\left(\bar{\psi}\gamma^{\mu}D_{\mu}\psi-m_{\psi}\bar{\psi}\psi\right)+\right.}$ (10)
$\left.+\frac{i}{2}\eta_{\mu\nu}\partial^{\rho}\bar{\psi}\gamma_{\rho}\psi\right]+\left[\frac{1}{4}\eta_{\mu\nu}F^{\lambda\rho}F_{\lambda\rho}-F_{\mu\lambda}F_{\nu}^{\lambda}\right]+\left[\eta_{\mu\nu}D^{\rho}H^{\dagger}D_{\rho}H+\eta_{\mu\nu}V(H)+\right$.
$\left.+D_{\mu}H^{\dagger}D_{\nu}H+D_{\nu}H^{\dagger}D_{\mu}H\right],$
and
$T_{\mu\nu}^{DM}=\left(\partial_{\mu}S\right)\left(\partial_{\nu}S\right)-\frac{1}{2}\eta_{\mu\nu}\left(\partial^{\rho}S\right)\left(\partial_{\rho}S\right)+\frac{1}{2}\eta_{\mu\nu}m_{S}^{2}S^{2},$ (11)
$T^{\mu}_{\mu}=\Sigma_{f}m_{f}\overline{f}f-2m^{2}_{W}W^{+}_{\mu}W^{-\mu}-m^{2}_{Z}Z_{\mu}Z^{\mu}+(2m^{2}_{h_{0}}h_{0}^{2}-\partial_{\mu}h_{0}\partial^{\mu}h_{0})+$... (12)
The states that diagonalize the kinetic energy and have canonical normalization $h$ and $\phi$ are given by [4]
$\left(\begin{array}[]{c}{h_{0}}\\ {\phi_{0}}\end{array}\right)=\left(\begin{array}[]{cc}{1}&{6\xi\gamma/Z}\\ {0}&{-1/Z}\end{array}\right)\left(\begin{array}[]{cc}{\cos\theta}&{\sin\theta}\\ {-\sin\theta}&{\cos\theta}\end{array}\right)\left(\begin{array}[]{c}{h}\\ {\phi}\end{array}\right)=\left(\begin{array}[]{cc}{d}&{c}\\ {b}&{a}\end{array}\right)\left(\begin{array}[]{c}{h}\\ {\phi}\end{array}\right),$ (13)
where $Z^{2}=1+6\gamma^{2}\xi\left(1-\,\,6\xi\right)=\beta-36\xi^{2}\gamma^{2}$ is the coefficient of the radion kinetic term after undoing the kinetic mixing, $\gamma=\upsilon/\Lambda_{\phi},\upsilon=246$ GeV, $a=-\dfrac{cos\theta}{Z},b=\dfrac{sin\theta}{Z},c=sin\theta+\dfrac{6\xi\gamma}{Z}cos\theta,d=cos\theta-\dfrac{6\xi\gamma}{Z}sin\theta$. The parameters a, b, c, d define the mixing between the $\xi=0$ states and the $\xi\neq 0$ mass eigenstates [11].
    The mixing angle $\theta$ is
$\tan 2{\theta}=12{\gamma\xi Z}\frac{m_{h_{0}}^{2}}{m_{h_{0}}^{2}\left(Z^{2}-36\xi^{2}\gamma^{2}\right)-m_{\phi_{0}}^{2}},$ (14)
where $m_{h_{0}}$ and $m_{\phi_{0}}$ are the Higgs and radion masses before mixing. The new physical fields h and $\phi$ in (13) are Higgs-dominated state and radion. The square masses are given by
$m_{h,\phi}^{2}=\frac{1}{2Z^{2}}\left[m_{\phi_{0}}^{2}+\beta m_{h_{0}}^{2}\pm\sqrt{(m_{\phi_{0}}^{2}+\beta m_{h_{0}}^{2})^{2}-4Z^{2}m_{\phi_{0}}^{2}m_{h_{0}}^{2}}\right].$ (15)
There are four independent parameters $\Lambda_{\phi},\,\,m_{h},\,\,m_{\phi},\,\,\xi$ that must be specified to fix the state mixing parameters. We consider the case of $\Lambda_{\phi}=5$ TeV and $\frac{m_{0}}{M_{P}}=0.1$ which makes the radion stabilization model most natural [6].
     A compelling question arises regarding whether a intuitive and tractable alternative exists for modeling unparticles. A natural theoretical candidate for this purpose is found in models based on warped extra dimensions. This idea is mentioned in RS model [28, 29]. We use couplings of the SM particles with the Higgs, radion and unparticle listed in detail in Ref.[5, 19, 17, 14] for evaluating in the next section.
@@ II Theoretical framework of the Higgs - radion mixing in Randall-Sundrum model
III The influence of the scalar unparticle and polarization on the exclusive W boson hadronic decays in the final state at muon colliders
@@
Influence of unparticle and polarization on properties of high energy colliders have been intensively studied in Refs.[20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 8]. The production of dark matter and unparticles associated with Z boson at 8 TeV via pp collisions has been analyzed in Ref.[31]. Recently, the scalar unparticle signals at LHC 14 TeV are investigated in Ref.[32]. Phenomenology of heavy neutral gauge boson at muon colliders are presented in detail in [49] which indicate that it can provide an unprecedented potential in probing new physics beyond the SM.
@@
     Study for the exclusive boson hadronic decay with the ATLAS detector at 13 TeV is considered in detail in Ref.[48]. Recently, mass biases in reconstructing exclusive radiative hadronic decays of W boson at the LHC are given in Ref.[50]. While rare and exclusive few-body decays of the Higgs, Z and W bosons and the top quark are studied in Ref.[51], the W boson’s exotic decay channel is concerned in Ref.[52]. However, the influence of the unparticle and polarization at muon colliders through hadronic decays of W boson has not yet been invested. In this section, by using the rare decay channels of $W^{\pm}\rightarrow\pi^{\pm}\gamma$, $W^{\pm}\rightarrow K^{\pm}\gamma$, $W^{\pm}\rightarrow\rho^{\pm}\gamma$, we investigate the influence of the scalar unparticle and polarization on the hardronic production at muon colliders in the RS model.
Figure 1: Feynman diagrams for $\mu^{+}\mu^{-}\rightarrow W^{+}W^{-}\rightarrow\pi^{-}\pi^{+}\gamma\gamma/K^{-}K^{+}\gamma\gamma/\rho^{-}\rho^{+}\gamma\gamma$ collisions. $M^{\pm}$ stands for the $\pi^{\pm},K^{\pm},\rho^{\pm}$ .
@@
$M_{Z}=-i\dfrac{g}{2\cos\theta_{W}(q^{2}_{s}-m^{2}_{Z})}\overline{v}(p_{2})\gamma^{\sigma}\left(-\frac{1}{2}+2\sin^{2}\theta_{W}-\frac{1}{2}\gamma_{5}\right)u(p_{1})\left(\eta_{\sigma\beta}-\dfrac{q_{s\sigma}q_{s\beta}}{m^{2}_{Z}}\right)\varepsilon^{*}_{\mu}(k_{1})\Gamma_{WWZ}^{\beta\mu\nu}\varepsilon^{*}_{\nu}(k_{2})$ (19)
@@
$M_{U}=i\overline{g}_{\mu\mu U}\overline{g}_{WWU}\dfrac{A_{d_{U}}}{2sin(d_{U}\pi)}|q^{2}_{s}|^{d_{U}-2}e^{-i\pi(d_{U}-2)}\overline{v}(p_{2})u(p_{1})\varepsilon^{*}_{\mu}(k_{1})\left[\left(k_{1}k_{2}\right)\eta^{\mu\nu}-k_{1}^{\nu}k_{2}^{\mu}\right]\varepsilon^{*}_{\nu}(k_{2}).$ (22)
Here, $q_{s}=p_{1}+p_{2}=k_{1}+k_{2}$, $s=(p_{1}+p_{2})^{2}$ is the square of the collision energy, $M_{U}$ is the contribution by the scalar unparticle, which is important for the described process. $\overline{g}_{\mu\mu\phi},\overline{g}_{\mu\mu h},\overline{g}_{W\phi},\overline{g}_{Wh}$ are the couplings of Higgs/radion shown in Ref.[5]. $\overline{g}_{\mu\mu U},\overline{g}_{WWU},A_{d_{U}}$ are given by Ref.[19, 14]. Anomalous couplings are shown in Ref.[53] as follows
@@
The parameters $\Delta\kappa_{V}$ and $\lambda_{V}(V=\gamma,Z)$ serve to quantify the strength of the beyond SM physics. Experiment results from W boson characterization constrain $\Delta\kappa_{V}$, indicating that these corrections cannot exceed a few percent. The parameter $\lambda_{V}$ can be identified as follows: $\lambda_{V}\rightarrow\lambda_{V}/(1+s/\Lambda^{2})^{2}$ [54].
@@
$M_{t}=-\frac{g^{2}}{2}\varepsilon^{*}_{\nu}(k_{2})\overline{v}(p_{2})\gamma^{\mu}P_{L}\frac{\widehat{q}_{t}}{q^{2}_{t}-m^{2}_{\nu}}\varepsilon^{*}_{\mu}(k_{1})\gamma^{\nu}P_{L}u(p_{1}).$ (26)
The total cross-section for the whole process can be calculated as follows
@@
where $\psi=(\overrightarrow{p}_{1},\overrightarrow{k}_{1})$ is the scattering angle. The model parameters are chosen as $\lambda_{\mu\mu}=\lambda_{WW}=\lambda_{0}=1$, $\xi=1/6$ [5], $\Lambda_{\phi}=5$ TeV [6], $m_{h}$ = 125 GeV, $m_{\phi}$ = 125 GeV [33]. The integrated luminosity scaling of a high energy muon collider reaches 10 $\text{ab}^{-1}$ at a center-of-mass energy of $\sqrt{s}=10$ TeV [40, 41]. The bounds on the anomalous $W^{-}W^{+}\gamma$ and $W^{-}W^{+}Z$ couplings are provided by the LEP, Tevatron and LHC experiments. The ATLAS collaboration has updated the best available constraints on anomalous couplings $\Delta k_{\gamma}$, $\lambda_{\gamma}$, $\Delta k_{Z}$, $\lambda_{Z}$ obtained as follows: $\Delta k_{\gamma}\in[-0.135,0.190]$, $\lambda_{\gamma}\in[-0.065,0.061]$, $\Delta k_{Z}\in[-0.061,0.093]$, $\lambda_{Z}\in[-0.062,0.065]$ [56]. From the formulas (28 - 30) above and branching ratio of W boson in the soft-collinear effective theory (SCET) and the light-cone distribution amplitude (LCDA) framework shown in Ref.[51], we give estimates in detail for the cross-sections as follows: